Write the equation of the line which passes through (0, -4) and is parallel to the line 3x - y = 6.


Answer by mananth(16946)

line equation 3x-y=6
put it in the form y = mx +b
y=3x-6
the slope is 3
and b = -6
the line parallel to this will have the same slope.
m=3
plug the value (0,-4) in the equation y= mx+b
b=-4
y=3x-4 is the line passing through (0,-4)